OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2

OpenStudy (anonymous):

y-y1=m(x-x1) y-0=-2(x-3) y= -2x+6 2x+y=6 is the required equation of the line. @glovesu

OpenStudy (anonymous):

can you explain how you did that? @dpasingh

OpenStudy (ranga):

y = mx + b m = -2. So y = -2x + b The line passes through (3,0). That is x = 3, y = 0. Plug these two numbers in y = -2x + b and solve for b. 0 = (-2)(3) + b. b = 6 So y = -2x + 6

OpenStudy (calculusxy):

Okay we will do the problem step by step so that you can understand it. Since I am using my iPad I can't use the equation format to make it look easier and in practical way. So there is this formulae conspiring of y=mx+b I will give you the main aspect or the keywords necessary that usually professors look for. m=slope b=y-intercept Let's now get started We will now substitute the numbers to the variables. 0=(-2)(3)+b Now solve for b. This is easy because -2*3=6. So now that use know, b=6.